OpenStudy (anonymous):

John rolls a number cube twice. What is the probability that the sum of the 2 rolls is less than 7, given that the first roll is a 1? @aum @IMStuck @TinkerbellGirl @toxicsugar22 @TaylorS11 @TheKavMan @Heart_Broken_Kid @aaronq @abb0t @AnnalieseBronz @2345 @AngelWilliams16 @aegidious

OpenStudy (anonymous):

Sorry, but I don't understand probability at all :/ I can do hard facts, but not chance

OpenStudy (cj49):

1+1=2 1+2=3 1+3=4 1+4=5 1+5=6 1+6=7 5/6

OpenStudy (anonymous):

Thanks